OpenStudy (anonymous):

@shamil98 can you help me with this one?? please

OpenStudy (anonymous):

\[\sqrt[5]{8m^2n^4}*\sqrt[5]{20m^4n}\]

OpenStudy (anonymous):

\[\sqrt[5]{160m^8n^4}\]

OpenStudy (anonymous):

is that the next step

OpenStudy (shamil98):

Hmm.

OpenStudy (anonymous):

????

OpenStudy (shamil98):

When you multiply exponents you add them..

OpenStudy (shamil98):

\[\sqrt[5]{8m^2n^4 \times 20m^4 n}\]

OpenStudy (anonymous):

ohhhhhhhhhhhh

OpenStudy (shamil98):

\[\sqrt[5]{160m^6n^5}\]

OpenStudy (anonymous):

and them simplify right???

OpenStudy (shamil98):

Yes.

OpenStudy (shamil98):

You can only get the 5th root of 160 in this though, so what would that be?

OpenStudy (anonymous):

16mn squareroot 2m

OpenStudy (anonymous):

is that right??

OpenStudy (shamil98):

no, the m^6 and n^5 stay in \[\sqrt[5]{m^6 n^5}\]

OpenStudy (shamil98):

all you have to do is find \[\sqrt[5]{160}\]

OpenStudy (anonymous):

but i thought in simplest radical form you do the exponents as well as the numbers

OpenStudy (shamil98):

well actually \[\sqrt[5]{n^5} = n\]

OpenStudy (shamil98):

\[\sqrt[n]{a^m} \] = a^m/n

OpenStudy (shamil98):

Luigi feel free too butt in if i'm doing this wrong .-.

OpenStudy (anonymous):

\[16mn \sqrt{2m}\]

OpenStudy (anonymous):

where did luigi come form

OpenStudy (shamil98):

\[\sqrt[5]{160m^6 n^5} = 2n \sqrt[5]{5m^6}\]

OpenStudy (shamil98):

right?

OpenStudy (anonymous):

oh you meant the guy luigi. i thought video game luigi

OpenStudy (luigi0210):

Sham has the right idea so far. But \[\LARGE \sqrt[5]{m^6}=m \sqrt[5]{m}\]

OpenStudy (shamil98):

Oh.

OpenStudy (luigi0210):

Everything else is fine, nice job.

OpenStudy (shamil98):

\[2nm \sqrt[5]{5m}\] Then.

OpenStudy (anonymous):

yup thanks luigi

OpenStudy (luigi0210):

Anytime buds.